# DiffuMatting: Synthesizing Arbitrary Objects with Matting-level Annotation

### v1

@@ Abstract
Due to the difficulty and labor-consuming nature of getting highly accurate or matting annotations, there only exists a limited amount of highly accurate labels available to the public. To tackle this challenge, we propose a DiffuMatting which inherits the strong Everything generation ability of diffusion and endows the power of ‘matting anything’. Our DiffuMatting can 1). act as an anything matting factory with high accurate annotations 2). be well-compatible with community LoRAs or various conditional control approaches to achieve the community-friendly art design and controllable generation. Specifically, inspired by green-screen-matting, we aim to teach the diffusion model to paint on a fixed green screen canvas. To this end, a large-scale green-screen dataset (Green100K) is collected as a training dataset for DiffuMatting. Secondly, a green background control loss is proposed to keep the drawing board as a pure green color to distinguish the foreground and background. To ensure the synthesized object has more edge details, a detailed-enhancement of transition boundary loss is proposed as a guideline to generate objects with more complicated edge structures. Aiming to simultaneously generate the object and its matting annotation, we build a matting head to make a green-color removal in the latent space of the VAE decoder. Our DiffuMatting shows several potential applications (e.g., matting-data generator, community-friendly art design and controllable generation). As a matting-data generator, DiffuMatting synthesizes general object and portrait matting sets, effectively reducing the relative MSE error by 15.4% in General Object Matting and 11.4% in Portrait Matting tasks.
@@ 1 Introduction
  • As a role of matting-data factor, we synthesize the general object and portrait matting set respectively and demonstrate that the inclusion of synthesized data leads to a reduction of 15.4% relative MSE error in General Object Matting and 11.4% relative MSE error in Portrait Matting task.
@@ 3 Methodology
Motivation and Applications. Despite the widespread use of large-scale models for image generation, there is little research focus on matting-level transparent image generation. Incorporating transparency is extensively used in the majority of visual content editing software and workflows to facilitate the composition and creation of content, which is significantly needed by the commercial market. In addition, the scarcity of matting-level data hinders the development of matting algorithms towards high accuracy and strong robustness. Driven by these demands, we propose DiffuMatting to generate the 4-channels images with the high-accurate matting channels for the following core-meaning applications: 1). Data factory. DiffuMatting can act as a data factory for downstream matting tasks in Fig. 6 and 7. 2). Community-friendly art design and controllable generation. DiffuMatting exhibits excellent compatibility with community LoRAs and various conditional control approaches in Fig. 10 and Fig. 11. It also supports the image composition in desired scenario in Fig. 8.
@@ 3.2 Preliminaries of Cross-attention Mechanism
where a Variational AutoEncoder $\mathcal{E}$ compress input image $x$ into low-dimension latent space $z$. A conditional U-Net denoiser $\epsilon_{\theta}$ aims to estimate noise $\theta$ in latent space with the aid of timestep $t$, $t$-th noisy latent representation $z_{t}$, and other text-prompt conditions $C$ extracted by text-encoder. The cross-attention mechanism is implemented to fuse the visual and textual embedding via:
@@ 4.3 Matting-level Annotation Analysis
Figure 7: The portrait matting results using our synthesis 10K Portrait-matting set based on the P3M-Net [20].
@@ 4.4 Applications
Downstream Matting Task. To verify the benefits of our matting-level annotation to the down-streaming matting-level task, we analyze our synthesized images on two mainstream matting tasks (General-object and Portrait matting). 1). General Object Matting: We synthesize the 10K general-object matting set (GOM) based on the around 500 object classes guided by the detailed captions of Green100K dataset. We train a general matting model based on the commonly-used Indexnet [26] under the same default setting on various training sets, consisting of images from the general-object matting set (GOM) and Deep Image Matting (DIM) dataset. The synthetic to real data ratio $\gamma$ denotes that $\gamma$ ratio of images from the DIM dataset and ($1-\gamma$) ratio of images from the synthesized GOM. As shown in Fig. 6, the combined dataset (highlighted in bold) used for matting gets better results (15.4% relative error lower on MSE, 9.4% on CONN) than only using the DIM dataset (Ratio $\gamma$=1), indicating that synthesized images by our DiffuMatting is useful for downstream general object matting even though it stills exists domain difference between the real and synthesized dataset.
@@
2). Portrait Matting: We also generate the 10K human-portrait-matting data (PM) including various age stages, races, genders, and nationalities. The largest public portrait dataset available is the P3M dataset with 10K images. The quantitative results are shown in Fig. 7. Adding our synthesis 10K portrait-matting set consistently improves the performance of network, and the optimal ratio is $4/5$ meaning that $1/5$ from the synthesized PM set and $4/5$ from the real P3M dataset. The ratio setting decreases the MSE error by 11.4% than the results of only using the real set.

### v2

@@ Abstract
Due to the difficulty and labor-consuming nature of getting highly accurate or matting annotations, there only exists a limited amount of highly accurate labels available to the public. To tackle this challenge, we propose a DiffuMatting which inherits the strong Everything generation ability of diffusion and endows the power of ‘matting anything’. Our DiffuMatting can 1). act as an anything matting factory with high accurate annotations 2). be well-compatible with community LoRAs or various conditional control approaches to achieve the community-friendly art design and controllable generation. Specifically, inspired by green-screen-matting, we aim to teach the diffusion model to paint on a fixed green screen canvas. To this end, a large-scale green-screen dataset (Green100K) is collected as a training dataset for DiffuMatting. Secondly, a green background control loss is proposed to keep the drawing board as a pure green color to distinguish the foreground and background. To ensure the synthesized object has more edge details, a detailed-enhancement of transition boundary loss is proposed as a guideline to generate objects with more complicated edge structures. Aiming to simultaneously generate the object and its matting annotation, we build a matting head to make a green-color removal in the latent space of the VAE decoder. Our DiffuMatting shows several potential applications (e.g., matting-data generator, community-friendly art design and controllable generation). As a matting-data generator, DiffuMatting synthesizes general object and portrait matting sets, effectively reducing the relative MSE error by 15.4% in General Object Matting. The dataset is released in our project page at https://diffumatting.github.io.
@@ 1 Introduction
  • As a role of matting-data factor, we synthesize the general object and portrait matting set respectively and demonstrate that the inclusion of synthesized data leads to a reduction of 15.4% relative MSE error in General Object Matting.
@@ 3 Methodology
Motivation and Applications. Despite the widespread use of large-scale models for image generation, there is little research focus on matting-level transparent image generation. Incorporating transparency is extensively used in the majority of visual content editing software and workflows to facilitate the composition and creation of content, which is significantly needed by the commercial market. In addition, the scarcity of matting-level data hinders the development of matting algorithms towards high accuracy and strong robustness. Driven by these demands, we propose DiffuMatting to generate the 4-channels images with the high-accurate matting channels for the following core-meaning applications: 1). Data factory. DiffuMatting can act as a data factory for downstream matting tasks in Fig. 6. 2). Community-friendly art design and controllable generation. DiffuMatting exhibits excellent compatibility with community LoRAs and various conditional control approaches in Fig. 9 and Fig. 10. It also supports the image composition in desired scenario in Fig. 7.
@@ 3.2 Preliminaries of Cross-attention Mechanism
where a Variational AutoEncoder $\mathcal{E}$ compress input image $x$ into low-dimension latent space $z$. A conditional U-Net denoiser $\epsilon_{\theta}$ aims to estimate noise $\theta$ in latent space with the aid of timestep $t$, $t$-th noisy latent representation $z_{t}$, and other text-prompt conditions $C$ extracted by text-encoder.
The cross-attention mechanism is implemented to fuse the visual and textual embedding via:
@@ 4.4 Applications
Downstream Matting Task. To verify the benefits of our matting-level annotation to the down-streaming matting-level task, we analyze our synthesized images on General-object matting. 1). General Object Matting: We synthesize the 10K general-object matting set (GOM) based on the around 500 object classes guided by the detailed captions of Green100K dataset. We train a general matting model based on the commonly-used Indexnet [26] under the same default setting on various training sets, consisting of images from the general-object matting set (GOM) and Deep Image Matting (DIM) dataset. The synthetic to real data ratio $\gamma$ denotes that $\gamma$ ratio of images from the DIM dataset and ($1-\gamma$) ratio of images from the synthesized GOM. As shown in Fig. 6, the combined dataset (highlighted in bold) used for matting gets better results (15.4% relative error lower on MSE, 9.4% on CONN) than only using the DIM dataset (Ratio $\gamma$=1), indicating that synthesized images by our DiffuMatting is useful for downstream general object matting even though it stills exists domain difference between the real and synthesized dataset.
@@ 5 Conclusion
Acknowledgments. This work was supported by National Key R&D Program of China (No. 2022ZD0118202), in part by the National Natural Science Foundation of China (No. 62072386), in part by Yunnan Provincial Major Science and Technology Special Plan Project (No. 202402AD080001), in part by by Henan Province key research and development project (No. 231111212000) and the Open Foundation of Henan Key Laboratory of General Aviation Technology (No. ZHKF-230212).